OpenStudy (anonymous):

What is the geometric mean of 5 and 15? 75 37.5

OpenStudy (anonymous):

75 37.5

hartnn (hartnn):

geometric mean of a and b is \(\sqrt{ab}.\)

hartnn (hartnn):

put a=5, b= 15

OpenStudy (anonymous):

so 5 X 15 then square root it?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

so square root of 60?

hartnn (hartnn):

15*5 = 75

hartnn (hartnn):

so , geometric mean = \(\sqrt {75}\)

OpenStudy (anonymous):

lol my bad but how do you simplify that

hartnn (hartnn):

isn't \(\sqrt {75}\) in options ?

hartnn (hartnn):

brooke, thats arithmetic mean.

hartnn (hartnn):

anyways, \(\sqrt{75}=\sqrt{3*25}=\sqrt{25}\sqrt{3}=5\sqrt{3}\)

OpenStudy (anonymous):

ya the options are 5 square root 3 75 3 square root 5 37.5

OpenStudy (anonymous):

ok so first one thank you

hartnn (hartnn):

u got how ?

OpenStudy (anonymous):

ya thank you i just needed help on first part but you showed me how thank you very much

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]